# A Compact, Low-cost Force and Torque Sensor for Robot Fingers with LED-based Displacement Sensing

### v1

@@ Abstract
Force/torque sensing is an important modality for robotic manipulation, but commodity solutions, generally developed with other applications in mind, do not generally fit the needs of robot hands. This paper introduces a novel method for six-axis force/torque sensing, using LEDs to sense the displacement between two plates connected by a transparent elastomer. Our method allows for finger-size packaging with no amplification electronics, low cost manufacturing, and easy integration into a complete hand. On test forces between 0-2 N, our prototype sensor exhibits a mean error between 0.05 and 0.07 N across the three force directions, suggesting future applicability to fine manipulation tasks.
@@
Force/Torque Sensing, Robot Hand, Robotic Manipulation
@@ I Introduction
Force and torque (F/T) measurements are useful for dexterous manipulation [1, 2, 3, 4, 5], as integrating an F/T sensor into the fingers can either replace or complement tactile sensors. While tactile sensors convey either the exact location of contact or a pressure map, often with some information about contact force and area, an F/T sensor only provides a net wrench. In an ideal world, tactile sensors could thus render F/T sensors redundant, but, as tactile sensors offering full fingertip coverage are still not ubiquitous technology, F/T sensors can still have a role to play in robot fingers.
@@
However, F/T sensors are primarily developed with industrial applications in mind [6, 7, 8], leading to limitations when being used for dexterous manipulation. Manipulation often benefits from compliance [9, 10, 11, 12], whereas most F/T sensors are optimized to be as stiff as possible, thus requiring precision engineering and leading to a high cost. Industrial applications also tend to favor a higher resolution than is needed for manipulation, but compromise by limiting the overload protection [6]. This leads to sensors that are easily damaged in the uncertain environments that robotics need to interact with. Finally, off-the-shelf F/T sensors can be difficult to package and integrate into the hand [13, 14, 15]. Others have attempted to solve this problem by designing sensors for manipulations to be placed within the fingertips [16, 17, 18, 19]. This limits the design of the finger to the needs of the sensor, and can also limit the inclusion of other modalities within the finger.
@@
We thus set out to build an F/T sensor that addresses the needs of robotic manipulation. First, such a sensor must be easy to manufacture using low-cost components and processes, so that it is affordable and practical to include multiple in a multi-fingered hand. Second, it must have the size and profile suited for the base of a fingertip, which will increase ease of integration into a hand. This also frees up the finger design for alternate sensors, or allows more flexibility in design of a finger for manipulation. Third, some amount of compliance could be beneficial to manipulation, and thus there is no need to avoid compliance at all cost. Finally, the ideal sensor would have a Signal to Noise Ratio (SNR) that allows it to be useful for manipulation without sacrificing overload protection.
@@
Figure 1: Our force/torque sensor, standalone (left) and mounted to a fixed base and robot finger (right). We use LEDs mounted on two plates connected by a transparent elastomer to sense displacement between the plates and relate it to applied forces and torques. This method allows for a compact, fully integrated package at a low manufacturing cost.
@@
To that end, we develop a novel F/T sensor based on light transport. Our sensor consists of two plates, connected by an elastomeric layer allowing 6 Degrees of Freedom (DOF) between them. To measure this displacement, we use a network of LEDs on each plate as both light emitters and light receivers. As the plates shift relative to each other due to an externally applied wrench, the signal measured by the receivers changes due to the relative movement and positioning of the emitters. This allows us to track the displacement, which can be used as a proxy for 6-axis F/T between the two plates. We map the change in signal from the light receivers to F/T signal using a neural network. Overall, our primary contributions are:
  • To the best of our knowledge, we are introducing a novel modality for F/T sensing, using LEDs to track the displacement between plates and mapping the signal to force and torque.
  • Our method results in a compact, low-cost F/T sensor that is compliant, easy to integrate in robot fingers, and has an observed resolution of .06 N and 2.6 N-mm.
@@
TABLE I: Comparison of novel 6-axis F/T sensors developed in the context of manipulation
II Related Work
In Table 1, we organize other low-cost F/T sensors developed to address the needs in manipulation. For example, Guggenheim et al. embedded barometers in a rubber flexure and placed them between two rigid plates [20]. These plates are relatively large with the length and width being 75mm. The size is constrained by the number of barometer modules needed to get full 6-axis forces and torques.
Camera-based sensors are more compact as they only need a single camera module, but require a completely translucent or hollow flexure. Some low-cost F/T sensors use small cameras to track the displacement of a marker fixed on some flexure. Ouyang et al. suspended a fiducial marker above a small camera module, using a compression spring as the flexure [21]. However, this design requires a relatively tall structure at 51mm. The Visiflex, a camera-based F/T sensor developed by Fernandez et al., is much more compact [18]. However, it only tracks forces and torques applied directly on the surface of the sensor. Therefore, contact anywhere else on the manipulator will not be sensed. In contrast, Kim et al. developed a very compact and capable F/T sensor using cylindrical capacitors [19]. The overall package of the sensor is very small, especially compared to the other low-cost sensors listed in Table I. While force range is high, torque range is limiting, as is the absolute accuracy reported.
Overall, we note that different sensors in Table I shine in different respects. In this context, the method introduced in this paper allows for compact size (27mm diameter and 20mm height), high bandwidth (2.5KHz) and high absolute accuracy, at a low manufacturing cost (\$100 in small quantities). We have not yet investigated maximum forces and overload protection, as these are highly depended on the geometry of the sensor and can thus be further optimized. Overall, we believe each of these technologies holds promise for advancing manipulation, and different sensors will suit different use cases and application scenarios.
III Sensor Concept
F/T sensors, at a high level, measure displacement between two rigid plates. This displacement can be measured through a variety of different electronic modalities. The rigid plates are fixed with some electronics that allow us to measure very small displacements and map them back to the force and torques applied. These plates are then attached to each other by some flexure.
As shown in Fig. 2, the electronics fixed to each rigid plates in our sensor will be a circuit boards with receiving and emitting LEDs. The plates are then suspended in a solid cylinder of PDMS that will allow for elastic displacements. At a neutral position, the emitter LED and the receiving LED are directly above each other. When an external force is applied, the plates will displace relative to each other and the cone of the emitting LED will change how it shines on the receiving LED. This will yield a change in signal resulting from a displacement, which was induced by some force. We will then use a data-driven approach to map this change in LED signal to the force applied.
Figure 2: A diagram of our LED-to-LED sensor concept. We align two PCBs with LEDs directly above each other. We attach these circuit boards to rigid plates, we then suspend both of these plates in a translucent, elastic material. Assuming the bottom plate is fixed, a force applied to the top plate will displace the plate and change how to emitter LED shines on the receiving LED. The displacement will change the receiving signal, which we then will use in a data-driven approach to relate back to the force applied.
PDMS enables us to use light as a sensing modality due to its low attenuation. Light offers three primary advantages. First, it provides a fast response time, as photodiode readings are gated by the ADC rather than the diodes. Second, its components require minimal space, which aids with reducing our footprint. Third, we can select LEDs in the IR spectrum, making the sensor resistant to the majority of environment noise that light based sensors suffer from without the need to design a cover or skin.
Using LEDs as both a receiver and emitter gives us several advantages over the other modalities menioned in Table I. Photodiodes have both a photoconductive mode (operating the diode via reverse-biased) which provides higher response time and a photovoltaic mode (operating the diode with zero-bias) which minimizes noise due to dark current. Since photovoltaic mode is well within the response range and time we need, we opt for that in order to minimize the noise and help maximize SNR. The decision between LEDs and Photodiodes was made based on 3 factors. 1) LEDs have a smaller form factor than most photodiodes in the price range. 2) LEDs are optimized for operation in their wavelength range, minimizing noise from external sources. 3) LEDs have a smaller viewing angle, which is useful for increasing sensitivity given our size constraints.
IV Analysis of Single Emitter-Receiver Pair
Figure 3: A two-axis, servo-driven testing rig we used to verify that an LED emitter-receiver pair will give us sufficient sensitivity to small displacements. The testing rig allows for both vertical and horizontal displacement. The vertical axis is driven by lead screws actuated by a Ax-12A. The horizontal axis is driven by a lead-screw and actuated by an XM-430A. This testing rig allows us to measure signal changes at .1mm displacements.
We started with tests to verify both the validity of LEDs as a basic building block, and the assumptions that LEDs would be better suited for our purposes than photodiodes. We created simple test boards with single LEDs as emitters, single LEDs as receivers, and photodiodes as receivers. We attached them to a servo-driven testing rig, shown in Fig. 3, which allowed us to measure the signal change over a horizontal and vertical displacement. By using lead screws and precise servos, we were able to test displacements as small as 0.1mm. This process revealed that the LED-to-LED pairing was sufficiently sensitive, especially to small displacements. In fact, both its sensitivity and its Signal to Noise Ratio (SNR) was higher than the photodiode, with a much sharper detection cone.
@@ IV Analysis of Single Emitter-Receiver Pair
Figure 4: Graphs showing the sensitivity of LEDs and Photodiodes when used as receivers. Left: This graph shows the sharper cone response of the LED when compared to the photodiode. Right: This graph is a zoom in on the ADC signal received across a .1mm window. It shows the clearer SNR when using the LEDs as receivers versus the dedicated photodiodes. Because the signal is spread over a smaller area, there are sharper changes at small displacements, including in a .1 mm window.
To understand the effect PDMS would have on our signal, we needed a different approach, since the LEDs could not be freely displaced as with air. Instead, for our receiver board, we used an array of LEDs mounted every 2.5mm, and measured how the cone shifted from air to PDMS. Since PDMS has a larger index of refraction relative to air, our assumption under Snell’s Law was that this would lead to a smaller cone with a sharper signal concentrated on a smaller area. What we found is that not only was the signal decreasing due to attenuation, we also measured a wider emission cone than in air. This informed our decision to switch LEDs to one with an even narrower cone, so that our cone would be in the desired width in PDMS. Overall, our LED-to-LED testing revealed that using LEDs as both emitters and receivers leads to a high signal-to-noise ratio that is capable of distinguishing displacement below .1mm and will be a good choice for our F/T sensor.
V Complete Sensor Design
@@
Figure 5: Our F/T sensor is composed of two custom PCBs that have both LED emitters and clusters of LED receivers. When the top and bottom PCB boards are aligned directly above each other, our sensor has 3 total emitter-receiver pars. The PCBs are fastened to rigid, 3-D printed plates. The plates are then fixed above each other in a mold that we then fill with PDMS. Once cured, the two PCBs are suspended above each other in a solid cylinder of 10:1 PDMS. We use a single FFC connector to communicate between the top and bottom board. We use an additional FFC connector on the bottom board to communicate with external electronics. The rigid plates are fitted with threaded holes to mount external hardware.
@@
With the core concept verified, we implement a preliminary design for the sensor. As we wanted to mount this at the base of a finger, we constrained our design to a 1 inch (25.4 mm) diameter. The goal was to maximize the amount of data signals within the space. The primary size constraint became the op-amps, which meant we could support a maximum of 12 signals per board. As shown in the left of Fig. 5, we divided the signals into clusters of 4 receivers, so that each board had 3 clusters, for a total of 24 signals broken across 12 clusters. We now must suspend the PCB boards above each other and attach them with an elastic flexure.
@@
We fasten the bottom and top PCB boards, as shown in Fig. 5, each to a rigid, 3-D printed plate. The plates each have three M2 inserts that we can mount external hardware too. We use these inserts to attach these rigid plates to a mold that we then fill with an uncured mixture of PDMS. We cure the filled mold at $70^{\circ}C$ for 8 hours, which gives us the sensor shown in Fig. 5 and 1. The overall dimensions of the sensor is 20.82mm in height and a diameter of 27mm, which is just above our target diameter.
Mechanically, our biggest concerns related to the stiffness of PDMS and sheer resistance. The PDMS needed to be compliant enough to displace at low forces, but stiff enough to survive the larger forces. We cured the PDMS at a 10:1 ratio of base to curing agent, as we observed this best achieved the desired sensor stiffness and surface finish. We observed that the PDMS does not adhere well to the PCB face, which makes the sensor vulnerable to shear failure. To address this issue, we treated the PCB boards with silicone primer to promote strong adhesion between the PDMS and PCB faces. We also added a lip to the rigid plates that helps prevent the PDMS from shearing off of the PCB faces.
@@
Our electronics also need to take into account concerns with integration in a robot hand. To that end, we want to minimize the number of signals transmitted between the board and the controller. Using an on-board ADC provides four advantages: 1) It allows for higher quality signals than utilizing the built-in ADCs on the microcontroller. 2) It allows us to use SPI protocol with only 7 wires. This also enables chaining of the top plate and bottom plate of the sensor using the SPI protocol. 3) It allows us to only transmit digital signals across the cables, minimizing transmission noise. 4) We maintain higher response time by avoiding external multiplexing.
Figure 6: Predicted vs ground truth Force and Torques across all 6 DoFs. Results presented from a test set with low accuracy on force but high on torque and a test set with high accuracy on force but low accuracy on torque for comparison. Top: Results from a contact with mean force error of .547N and torque error of 1.23Nmm. The prediction follows the shape tightly for Fx and Fy, but struggles at points with the magnitude. The torques meanwhile are much more accurate, following both shape and magnitude. Bottom: Force error of .0494N and torque error of 2.10Nmm. This prediction consistnetly undervalues the magnitude of both force and torques.
VI Experiments
VI-A Data Collection
To gather the dataset, we mounted a 3D-printer, finger on top of the sensor and contacted it from all around its circumference. The sensor was attached to an ATI SI-130-10 for ground truth labels. The dataset gathered for this paper focused exclusively on forces coming horizontally at the finger, rather than vertically at the finger tip. Each data file consists of three contacts within a 20 second span. It included a range of forces at each height to ensure that force and torque were decoupled. The dataset focused on forces primarily between 0-2N and torques between 0-.2Nm, but contained data expanding all the way up to 10N of force.

### v2

# Compact LED-Based Displacement Sensing for Robot Fingers

@@ Abstract
In this paper, we introduce a sensor designed for integration in robot fingers, where it can provide information on the displacements induced by external contact. Our sensor uses LEDs to sense the displacement between two plates connected by a transparent elastomer; when a force is applied to the finger, the elastomer displaces and the LED signals change. We show that using LEDs as both light emitters an receivers in this context provides high sensitivity, allowing such an emitter and receiver pairs to detect very small displacements. We characterize the standalone performance of the sensor by testing the ability of a supervised learning model to predict complete force and torque data from its raw signals, and obtain a mean error between 0.05 and 0.07 N across the three directions of force applied to the finger. Our method allows for finger-size packaging with no amplification electronics, low cost manufacturing, easy integration into a complete hand, and high overload shear forces and bending torques, suggesting future applicability to complete manipulation tasks.
@@
Contact Sensing, LED emitter and receiver, Robot Hand, Robotic Manipulation
@@ I Introduction
Detecting the net forces or torques acting on a robotic finger is useful for dexterous manipulation [1, 2, 3, 4, 5], with the potential to replace or complement tactile sensors. Tactile sensors convey significantly more information, such as the exact location of contact or a pressure map, from which net finger forces can also be extracted. However, tactile sensors offering full fingertip coverage are still not ubiquitous technology. In their absence, information on net forces and torques can still be useful in robot fingers engaged in manipulation tasks.
@@
The most direct method for sensing such forces and torques is to add an off-the-shelf six-axis force and torque (F/T) sensor into the base of the finger. However, such F/T sensors are primarily developed with industrial applications in mind [6, 7, 8], leading to limitations when being used for dexterous manipulation. Manipulation often benefits from compliance [9], whereas most F/T sensors are optimized to be as stiff as possible, thus requiring precision engineering and leading to a high cost. Industrial applications also tend to favor a higher resolution than is needed for manipulation, but compromise by limiting the overload protection [6]. This leads to sensors that are easily damaged in the uncertain environments that robotics need to interact with. Finally, off-the-shelf F/T sensors can be difficult to package and integrate into the hand. Others have attempted to solve this problem by designing sensors to be placed within the fingertips [10, 11, 12]. This limits the geometry of the finger and can also limit the inclusion of other modalities within the finger.
@@
In this paper, we present the design of a displacement sensor intended to provide information about the net forces acting on a robot finger in a manner that addresses the needs of robotic manipulation. First, such a sensor must be easy to manufacture using low-cost components and processes, so that it is affordable and practical to include multiple in a multi-fingered hand. Second, it must have the size and profile suited for the base of a fingertip, which will increase ease of integration into a hand. This also frees up the finger design for alternate sensors, or allows more flexibility in design of a finger. Third, some amount of compliance could be beneficial to manipulation, and thus there is no need to avoid compliance at all cost. Finally, the ideal sensor would have a Signal to Noise Ratio (SNR) that allows it to be useful for manipulation without sacrificing overload protection.
@@
Figure 1: Our LED-based sensor, standalone (left) and mounted to a fixed base and robot finger (right). We use LEDs mounted on two plates connected by a transparent elastomer to sense displacement between the plates and relate it to applied forces and torques. This method allows for a compact, fully integrated package at a low manufacturing cost.
@@
To that end, we develop our sensor based on light transport. Our design, shown in Fig. 1, consists of two plates, connected by an elastomer layer that allows for 6 Degrees of Freedom (DOF) between them. When mounted at the base of a robot finger, any contact force applied to the finger will cause a displacement between these plates. To measure this displacement, we use a network of LEDs on each plate as both light emitters and light receivers. As the plates shift relative to each other due to an externally applied forces, the signal measured by the receiver LEDs changes due to the relative movement and positioning of the emitter LEDs. We record these signals, and use them to extract information about the forces acting on the finger.
A key feature of our work is that using LEDs as both emitters and receivers confers very high signal-to-noise ratio, significantly higher than using photodiodes. Combined with the small LED form factor, this allows us to sense very small displacements in a compact package. Using an off-the-shelf F/T sensor as ground truth, we show that our sensor has an observed resolution of .06 N and 2.6 N-mm in a compact, low-cost package that is compliant and easy to integrate in robot fingers as it requires no external amplification. To the best of our knowledge, this is the first time that LEDs have been integrated into a 6-DOF flexure, allowing for full contact information in a compact, complaint robot fingers. Our sensor is designed to be integrated at the base of the finger, which allows for sensing coverage of the entire finger and the ability to integrate other sensors in the finger. We hope that this direction can facilitate truly ubiquitous sensing integrated in robot fingers, with future applications in manipulation.
Figure 2: A diagram of our LED-to-LED sensor concept. We align two PCBs with LEDs directly above each other. We attach these circuit boards to rigid plates, we then suspend both of these plates in a translucent, elastic material. Assuming the bottom plate is fixed, a force applied to the top plate will displace the plate and change how to emitter LED shines on the receiving LEDs. We can then map this change to the contact wrench on the sensor that induced the displacement.
II Related Work
Many tactile sensors use LEDs due to their compact packaging and wide range of emission cones [13, 14, 15, 16, 17]. The LEDs in these sensors are embedded in a translucent elastomer, so that their emission cone shines on a receiving device. This receiver is often a photodiode [14, 16] or camera [15, 17, 18], but can also be an LED [19]. Regardless of the receiving device, when a force is applied to the surface of these sensors, the elastomer deforms and the emitter and receiver will displace. Using either a data-driven approach or deriving an analytical relationship, the change in electrical signal from this deformation can then be mapped to the contact applied. These tactile sensors are not able to cover the full surface of a robot finger, and therefore do not provide full sensing coverage all possible contacts. To detect the full contact, others have tried several modalities including strain-based transducers [20], capacitance [21], optical [22], and magnetic [12]. Although most of these sensors are too large to integrate in a robot finger, there are a few novel F/T sensors that are built particularly for manipulation.
In Table 1, we organize other low-cost sensors developed specifically to address the needs of contact wrench detection in robot fingers. For example, Guggenheim et al. embedded barometers in a rubber flexure and placed them between two rigid plates [23]. These plates are relatively large with the length and width being 75mm. The size is constrained by the number of barometer modules needed to get 6-axis forces and torques. Camera-based sensors are more compact as they only need a single camera module, but require a completely translucent or hollow flexure. Some low-cost F/T sensors use small cameras to track the displacement of a marker fixed on some flexure. Ouyang et al. suspended a fiducial marker above a small camera module, using a compression spring as the flexure [24]. However, this design requires a relatively tall structure at 51mm. The Visiflex, a camera-based F/T sensor developed by Fernandez et al., is much more compact [10]. However, it only tracks forces and torques applied directly on the surface of the sensor. Therefore, contact anywhere else on the manipulator will not be sensed. In contrast, Kim et al. developed a very compact and capable F/T sensor using cylindrical capacitors [11]. The overall package of the sensor is very small, especially compared to the other low-cost sensors listed in Table I. While force range is high, torque range is limiting, as is the absolute accuracy reported.
Overall, we note that different sensors in Table I shine in different respects. In this context, the method introduced in this paper allows for compact size (27mm diameter and 20mm height), high bandwidth (2.5KHz) and good absolute accuracy, a low manufacturing cost (\$100 in small quantities), overload forces up to 110N, and overload torques up to 1.62 N*m. Overall, we believe each of these technologies holds promise for advancing manipulation, and different sensors will suit different use cases and application scenarios.
@@ II Related Work
TABLE I: Comparison of novel 6-axis sensors developed to extract full contact information in the context of dexterous robot manipulation
@@
Figure 3: Graphs showing the sensitivity of LEDs and Photodiodes when used as light receivers. Left: This graph shows the sharper cone response of the LED when compared to the photodiode. Right: This graph is a zoom in on the ADC signal received across a .1mm window. It shows the clearer SNR when using the LEDs as receivers versus the dedicated photodiodes. Because the signal is spread over a smaller area, there are sharper changes at small displacements, including in a .1 mm window.
III Sensor Design
III-A Concept
Our conceptual diagram of our sensor is shown in Fig. 2. It consists of two rigid plates connected by an elastomeric layer that allows 6-DOF displacements. The electronics fixed to each rigid plate consist of a circuit boards with emitter LEDs and clusters of receiving LEDs. At the neutral position, the emitter LED is directly above the center of the receiving cluster. When an external force is applied, the plates will displace relative to each other. This displacement will result in a change in signal for the receiver LEDs due to the relative motion of the corresponding emitter LED. We can then map this change in signal back to the contact wrench applied.
III-B Analysis of Single Emitter-Receiver Pair
Displacement transduction via light measurements offers three primary advantages. First, it provides a fast response time, as diode readings are gated by the ADC rather than the diodes. Second, its components require minimal space, which aids with reducing our footprint. Third, we can select LEDs in the IR spectrum, making the sensor resistant to the majority of external environment light. Adding an opaque wrapping to the outside of our sensor, we can further prevent any external light interference from external infrared light.
Using LEDs as both a receiver and emitter gives us several advantages over the other modalities mentioned in Table I. First, light can be sensed at very high frequencies with minimal electrical noise. There are several devices we could use as our receiver, but we opted to use LEDs as the receivers. In our case, the decision of using receiving LEDs over photodiodes was made based on 3 factors. 1) LEDs have a smaller form factor than most photodiodes in the price range. 2) LEDs are optimized for operation in their wavelength range, making it easier to design features to minimize external interference. 3) LEDs have a smaller viewing angle, which is useful for increasing sensitivity given our size constraints.
We first verify the validity of LEDs as both emitters and receivers, and compare to the alternative using photodiodes on the receiving end. To that end, we measured the response to relative horizontal displacements of a single emitter-received pair separated by an air gap, where the receiver consisted of either and LED or a photodiode. The results, plotting receiver signal vs. relative horizontal displacement of the emitter, are shown in Fig. 3. We observe that the LED receiver exhibits much higher sensitivity compared to its photodiode counterpart. In fact, the LED receiver can identify displacements as small as 0.01 mm, whereas the corresponding signal changes for a photodiode receiver are comfortably inside the noise level. Based on this result, we proceed with LED-LED pairs for the overall sensor design.
@@ III-B Analysis of Single Emitter-Receiver Pair
Figure 4: Our sensor is composed of two custom PCBs that have both LED emitters and clusters of LED receivers. When the top and bottom PCB boards are aligned facing each other, our sensor has 3 total emitter-receiver pars. The PCBs are fastened to rigid, 3-D printed plates. The plates are then fixed above each other in a mold that we then fill with PDMS and cure to get the sensor on the right. We use a single JST connector to communicate between the top and bottom board. We use an additional JST connector on the bottom board to communicate with external electronics. The rigid plates are fitted with threaded holes to mount external hardware. Our sensor is 27 mm in diameter, 20mm in height, and weighs 16.4g.
III-C Complete Sensor Design
@@
With the core concept of using identical LEDs as emitters and receivers verified, we implemented a preliminary design for the sensor. As we wanteded to mount this at the base of a finger, we constrained our design to a 1 inch (25.4 mm) diameter. The goal was to maximize the amount of data signals within the space. Based on our LED analysis, we used a compact, infrared LED (SFH 4056) with an emission cone of $20^{\circ}$. The primary size constraint became the op-amps, which meant we could support a maximum of 12 signals per board. As shown in the left of Fig. 4, we divided the signals into clusters of 4 receivers, so that each board had 3 clusters, for a total of 24 signals broken across 12 clusters. We then suspended the PCB boards above each other and attached them with an elastic flexure.
@@
We fastened the bottom and top PCB boards, as shown in Fig. 4, each to a rigid, 3-D printed plate. The plates each have three M2 inserts that we can mount external hardware too. We used these inserts to attach these rigid plates to a mold that we then filled with an uncured mixture of PDMS. We cured the filled mold at $70^{\circ}C$ for 8 hours, which gave us the sensor shown in Fig. 4. The overall dimensions of the sensor is 20.82mm in height and a diameter of 27mm, which is just above our target diameter. We then covered the sensor with a thin, opaque rubber shrink-wrap tube. This cover inhibits interference from external light sources, but also helps adds to the sensor robustness. We cured the PDMS at a 10:1 ratio of base to curing agent, as we observed this best achieved the desired sensor stiffness and surface finish. Moreover, this ratio produced the most consistent behavior across multiple sensors.
To increase PDMS adherence to the plates, we treated the PCB boards with silicone primer to promote strong adhesion between the PDMS and PCB faces. We also added a lip to the rigid plates that extends past the PCB surface to help the PDMS from shearing off the PCB. The rubber wrapping also increases the stiffness of our sensor and adds robustness under large displacements, especially in shear.
@@
Our electronics also need to take into account concerns with integration in a robot hand. To that end, we want to minimize the number of signals transmitted between the board and the controller. Using an on-board ADC provides four advantages: 1) It allows for higher quality signals than utilizing the built-in ADCs on the micro-controller. 2) It allows us to use SPI protocol with only 7 wires. This also enables chaining of the top plate and bottom plate of the sensor using the SPI protocol. 3) It allows us to only transmit digital signals across the cables, minimizing transmission noise. 4) We maintain higher response time by avoiding external multiplexing.
Figure 5: Mechanical hysteresis testing over three consecutive loading and unloading cycles. We fixed our sensor, with its rubber cover, on a reference F/T sensor and then used a linear probe to displace the top plate by 1 mm. Plotting the force and torque vs displacement for three consecutive cycles of pokes, we observe some hysteresis that is relatively consistent between cycles.
Figure 6: Mechanical overload forces and torques for our sensor. We built five sensors and then used an ADMET instron machine to apply a horizontal force until the sensor failed. By fixing it to the machine in different ways, we tested our sensor in five major failure modes - shear, bending, torsion, tension, compression. All tests were done with the opaque, shrink wrapped rubber cover on. Our sensor does not break under very high forces and torques, which is ideal for contact rich manipulation in uncertain robot environments.
Figure 7: Predicted vs. ground truth forces and torques across all 6 DoFs. Each row represents a test poke on our sensor. The contact prediction model demonstrates strong performance for forces and torques along/about the x and y-axes but struggles particularly with force prediction along the z-axis.
IV Mechanical Characterization
A core design objective of this sensor is robustness against large displacements induced by high contact forces. We tested our sensor to failure in five failure modes - shear, bending, torsion, tension, and compression. We note that, since our sensor is intended to be mounted at the base of a finger, as shown in Fig. 1, contact forces will mostly cause shear and bending of our sensor. Additionally, all tests on our sensor were done with a shrink wrapped neoprene cover, as shown on the finger in Fig. 1. This cover improved the robustness of our sensor, as it added an extra layer of protection to the PDMS flexure, especially for shear and tension forces.
The magnitude and direction of the forces and moments for each failure mode relative to our sensor are shown in Fig. 6. In particular, we note that our sensor withstands shear forces of up to 110 N and bending torques of up to 1.62 N*m. The overload forces for torsion and compression are also high. Our sensor is weakest in tension, which is not a major concern as a robot finger would likely not experience large tension forces while doing manipulation in uncertain environments.
PDMS is a hysteric material, so we expect our 6-DOF flexure to display some time-dependent non-linearity [13, 25]. In Fig. 5, we plot the results of cyclic loading and unloading tests for our sensor. We used a linear probe to horizontally displace the top plate of our sensor while it was fixed to a reference, off-the-shelf F/T sensor. Displacing the top plate by 1 mm three consecutive times, we observe hysteresis that is relatively consistent over the cycles. While hysteresis is a common effect for sensors that use PDMS as an elastic flexure [25], it is also problematic for mapping the electrical signals to the force applied, and often displays behavior that is unique to each sensor. In future work, we aim to investigate different clear elastomers for reducing this effect, potentially in combination with more complex learning architectures that can aim to learn this behavior and reduce in post processing.
V Performance
In this section, we aim to quantify the performance of our design as a standalone sensor. Ultimately, we believe that the true value of a sensor designed for robot fingers lies in the performance improvements it enables for complete manipulation. If manipulation is performed via learned policies, such policies are likely to use raw sensor data as input (in our case, raw readings from the receiver LEDs), decreasing the importance of any intermediate representations. However, training such a policy exceeds the scope of this paper, which focuses on the design of the sensor itself.
Therefore, in order to quantify the standalone performance of our sensor, we test the ability to infer ground truth wrenches applied to the robotic finger from the raw sensor data. To that end, we mount our sensor on a commercial ATI SI-130-10 to obtain ground truth force and torque measurements. We then trained a simple machine learning model to predict applied force and torque based on our sensor data. We describe the experimental approach and results in this section.
V-A Data Collection
To collect the dataset, we mounted a 3D-printed finger-like shape on top of our sensor, as shown on the right in Figure 1. We made contact with it from all directions along its circumference. Each data file contains three contact events recorded within a 15-second span. To ensure force and torque were decoupled, we applied a range of forces at different heights. The dataset primarily captures forces between 0–5 N and torques between 0–0.2 Nm along each axis, with a maximum force magnitude limited to under 10 N.